OpenStudy (anonymous):

The function y = x2 is transformed to y = (x + 4)2. Which statement is true about the transformed function? It is an even function. It is an odd function. It is neither an even nor an odd function. It is both an even and an odd function.

OpenStudy (anonymous):

\(y=x^2\) is even as it has only even exponents \(y=(x+4)^2=x^2+8x+16\) has both even and odd exponents, and so is neither even nor odd